# From multitype branching Brownian motions to branching Markov additive processes

### v1

@@ 1.4 Related works
Branching Brownian motion (BBM) and branching random walks (BRW) are canonical probabilistic models, with significant applications in statistical physics and population biology. These processes provide insights for understanding phenomena ranging from the extremes of log-correlated random fields and the structure of mean-field spin glasses to the dynamics described by diffusion-reaction equations. The classical case of a single particle type has been extensively studied. For a comprehensive treatment, we refer the reader to [ShiBRW] and [Bovier].
For branching Lévy process, there have been a recent increasing interests; see [Profeta24, HJRS25, HRS24, RSZ25], to list just a few. An extended model with infinite branching rate has been introduced by [BerMal19] and the martingale convergence results have been studied in [Bertoin2018, MalleinShi2023].
The theory of branching processes has been generalized to the multitype case. Kesten and Stigum [Kesten1966] established a key limit theorem for the discrete-time multitype branching processes, using the famous $L\log L$ moment condition for the reproduction law. Athreya [Athreya1968] proved this result for continuous-time settings. In the context of branching Brownian motion, Ren and Yang [Ren2014] studied irreducible multitype cases, offering a probabilistic proof for the existence, uniqueness and asymptotic behaviours of the corresponding travelling wave solutions. More recently, Hou et al. [Hou2024] proved that the extremal process of an irreducible multitype branching Brownian motion converges weakly to a cluster point process. For reducible cases, two-type branching Brownian motions have been investigated in [BelloumMallein2021] and [MaRen2024]. Note that, a common assumption in these works on multitype BBM has been that all particle types share the same underlying motion process. This is distinct from our setting, where the underlying processes differ among types.
A further generalization considers types in a general measurable space, which introduces significant new challenges, even without the presence of the spatial motion. There are significant new challenges. For the analysis of survival properties, we refer to [AndDuc25] for an $L\log L$ condition in the model with countably many types, and to [MaiTou25] for a comprehensive treatment of the case with uncountably many types.
@@ 1.5 Examples: two-type BBMs
To make connections with results in [BlaEtAl25, Champneys95], let us exam in detail a specific case of our model: a two-type branching Brownian motion. Particles have two types $\{1,2\}$. For $i=1,2$, a type $i$ particle undergoes a Brownian motion with drift $\mathtt{a}_{i}\in\mathbb{R}$ and variance $\sigma^{2}_{i}>0$; it branches at rate $\beta_{i}\geq 0$ into offspring of the same type at its current location, with offspring numbers distributed according to $\mu_{i}$. Furthermore, at rate $q_{1}>0$ (for type $1$) or $q_{2}>0$ (for type $2$), the particle switches to the other type without displacement. All particles evolve independently. Let $m_{i}=\sum_{k\geq 1}\mu_{i}(k)k$, $i=1,2$. Assume $\beta_{1}(m_{1}-1)+\beta_{2}(m_{2}-1)>0$ and $\sigma_{1}^{2}+\sigma_{2}^{2}>0$ to avoid degenerate cases.
@@
A connection with in [BlaEtAl25]
This formulation of two-type branching Brownian motions encompasses the on-off branching Brownian motion models proposed by [BlaEtAl25], where the two types are referred to as “active” and “dormant”. When $\beta_{2}=\sigma_{2}=0$, we obtain the Variant I model defined by [BlaEtAl25, (1.16)]; when $\beta_{2}=\sigma_{1}=0$, we obtain the variant II model in [BlaEtAl25, (1.19)].
In [BlaEtAl25], although not explicitly stated, the last formula on page 10 has assumed $\sum k^{2}\mu_{1}(k)<\infty$. This already implies $\sum_{k}(k\log^{2}k)\mu_{i}(k)<\infty$. Compared to [BlaEtAl25], our work provides several extensions:
  • For the linear speed of the leftmost particle, the result matches [BlaEtAl25, Theorem 1.9].
  • When $\theta>\theta^{*}$, we deduce the same results on the additive martingale [BlaEtAl25, Proposition 2.7] and travelling waves [BlaEtAl25, Theorem 1.13].
  • When $0<\theta<\theta^{*}$, we have the same results on the additive martingale [BlaEtAl25, Theorem 2.3 and Proposition 2.4] and travelling waves [BlaEtAl25, Theorem 1.10].
  • When $\theta=\theta^{*}$, we include new results on martingale convergence, and also the existence and uniqueness of the travelling wave solutions, answering the open questions in [BlaEtAl25].
A connection with [Champneys95]
The model in [Champneys95] is recovered by setting the drifts to zero ($a_{i}=0$) and requiring strictly positive branching rates ($\beta_{i}>0$ for $i=1,2$). In [Champneys95], Since the production law is given by $\mu_{1}(2)=\mu_{2}(2)=1$, it also satisfies $\sum_{k}(k\log^{2}k)\mu_{i}(k)<\infty$. Our results are consistent with [Champneys95]:
  • For the linear speed of the leftmost particle, the result matches the speed in [Champneys95, Theorem 1.41].
@@
  • When $0<\theta<\theta^{*}$, the $\mathcal{L}^{1}$-convergence is consistent with [Champneys95, Theorem 1.39]. We also deduce the same existence and uniqueness results in [Champneys95, Theorem 1.41].
@@ 1. Generalization to non-local branching with infinite branching rate
It is straightforward to extend our models to include non-local branching governed by a point process. We can further accommodate infinite branching rates as in [BerMal19], using similar approximation methods given there. This generalization would establish a connection with multitype growth-fragmentations [DaSilva23, DaSilvaPardo24].
@@ 2. Necessary conditions for the non-triviality of the derivative martingale limit
For the single-type BBM case, (1.14) has been proven to be necessary and sufficient for the critical derivative martingale to converge to a non-trivial limit in [YangRen11]; similar results are also known as the Aïdékon–Chen condition for single-type BRWs [Aidekon13, Chen15], and for branching Lévy processes with infinite branching rates [MalleinShi2023]. We conjecture that an analogous statement still holds for our multitype model, namely the condition (1.14) is both necessary and sufficient. Indeed, we believe that the methods from [MalleinShi2023] should be applicable to our model; however, a full proof would require a further study of the perpetual integral of a conditioned MAP.
@@ 3. Uniqueness of travelling waves
In Theorem 1.8, the assumption of spectrally negative jumps is used to establish uniqueness. We believe this assumption is redundant. Extending the result to processes with two-sided jumps requires a more refined analysis. We expect that the techniques from [AlsMal22], which first establishes asymptotics of the travelling waves and then use these to prove uniqueness, are applicable.
@@ 4. Finer study of the leftmost position and extremal process
It would be interesting to develop a refined analysis of the leftmost position, e.g. by studying the convergence of $\min_{u\in\mathcal{N}_{t}}X_{u}(t)+\frac{\lambda(\theta^{*})}{\theta^{*}}t$ as well as the extremal process. This is closely related to the convergence of FKPP solutions to travelling waves and the precise information on the front propagation. The spine decomposition in the current work allows the methods in [ShiBRW] to be adapted to the multitype setting.
@@ 5. Infinitely many types
A natural generalization is to consider more general type spaces, with countable or uncountable infinitely many types. Comparing with finitely many types, this can lead to significantly different behaviours; for example, local extinction of each type would no longer be equivalent to global extinction of the entire population [AndDuc25]. This framework naturally connects to heterogeneous models of spatially dependent branching and movement. Developing such an extension would require studies on general Markov additive processes, where the underlying modulating processes are general Markov processes, as opposed to Markov chains on a finite space; see [MaiTou25] for recent development.
@@ 2.1 Preliminaries on MAPs
This section provides the necessary preliminaries on Markov additive processes (MAPs) defined as in Definition 1.1. We refer to [Iva2011, Chapter 2] and [DLK17, Appendix] for detailed discussions on this topic.
Recall from Proposition 1.2 that a MAP $(\chi,\Theta)$ is associated with a family of Lévy processes $(\chi_{i})_{i\in\mathcal{I}}$ with respective Laplace exponents $(\phi_{i})_{i\in\mathcal{I}}$, a Markov chain with intensity matrix $\Theta$ and a family of random variables $(U_{ij})_{i,j\in\mathcal{I}}$ with Laplace transform $G=(G_{ij})_{i,j\in\mathcal{I}}$. The matrix exponent of the MAP $F(\alpha)$ is given by (1.7). Since $\Theta$ is irreducible, the entries of the matrix $e^{F(\alpha)t}$ are all strictly positive. From the Perron-Frobenius (PF) theory, $F(\alpha)$ admits a PF eigenvalue which is real and larger than the real part of any other eigenvalues; see for example [seneta06, Theorem 1.1]. Based on the PF eigenvalue, we have the following law of large numbers.
Proposition 2.1 ( [Iva2011, Propositions 2.13, 2.15 and Lemma 2.14]).
@@
The following lemma for the perpetual integral of a MAP will be used to prove Theorem 3.2. For more studies on the perpetual integrals of Lévy processes we refer to [BDK, KolSav19].
@@
In this proof we denote by $C_{n}>0$ suitable constants for $n\geq 1$. We first treat the case when $\chi$ is non-lattice. Using [DLK17, Theorem 27], we have
@@
Here $(R_{i,j}^{+})_{i,j\in\mathcal{I}}$ are the potential measures for the ascending ladder height process associated with $(\chi,\Theta)$, and $(R_{i,j}^{-})_{i,j\in\mathcal{I}}$ are those for the descending ladder height process; see [DLK17, Equation (27)] for more details. When $\chi$ is lattice, then each Lévy process $\chi_{i}$ is a compound Poisson process on the same lattice $\{r\mathbb{Z}\}$ for some $r>0$, and the analysis is similar to discrete random walks. In this case, the renewal measures can be expressed via the (strong or weak) ascending ladders, analogous to the renewal measure for discrete random walks. Let $(R_{i,j}^{>})_{i,j\in\mathcal{I}}$ denote the renewal measures for the strongly ascending ladder height process of $\chi$, and let $({R}^{\leq}_{i,j})_{i,j\in\mathcal{I}}$ denote the renewal measures for the weakly descending ladder height process. Specifically, we define the renewal measures as follows. Let $T_{0}^{>}:=0$ (resp. $T_{0}^{\leq}:=0$) and define recursively for $n\geq 0$ that $T_{n+1}^{>}:=\inf\{t>T_{n}^{>}:\chi(t)>\chi(T_{n}^{>})\}$ (resp. $T_{n+1}^{\leq}:=\inf\{t>T_{n}^{\leq}:\chi(t)\leq\chi(T_{n}^{>})\}$). Let $H_{n}^{>}:=\chi(T_{n}^{>})$ (resp. $H_{n}^{\leq}:=\chi(T_{n}^{\leq})$) be the strongly ascending (resp. weakly descending) ladder heights. Then we define, for integer $m\geq 0$ and $k\leq 0$,
@@
Then by similar arguments as in the proof of [DLK17, Theorem 27], we deduce a lattice analogue (c.f. [Spitzer, Page 209 P3] for the random walk case) of (2.2):
$\mathbb{E}_{x,i}\Bigg[\int_{\mathbb{R}_{+}}f(\chi(s),\Theta(s))\mathbbm{1}_{\left\{\tau_{0}>s\right\}}\mathrm{d}s\Bigg]$ $\leq C_{1}\sum_{k\in\mathcal{I}}\sum_{j\in\mathcal{I}}\sum_{n=0}^{\infty}\sum_{m=0}^{\lfloor x/r\rfloor}R_{i,j}^{>}(\{nr\}){R}_{k,j}^{\leq}(\{mr\})f(x+(n-m)r,k)$ (2.4)
@@
It follows from the Markov renewal theorem (see e.g. [Lalley84] or [Asmussen, Section VII.4]) that they are non-negative and there exists $A>0$  [^5] Note that it holds in fact for any $A>0$ for the non-lattice case; for the lattice case, take $A$ larger than the span. such that
$R_{i,j}^{+}([nA,(n+1)A)),R_{i,j}^{-}([nA,(n+1)A)),R_{i,j}^{>}([nA,(n+1)A)),{R}_{i,j}^{\leq}([nA,(n+1)A))\leq C_{2}A,\qquad\forall n\geq 0.$
@@
$\leq\sum_{n\geq 0}\widehat{f}(nA,k)C_{2}A\leq C_{4}+C_{2}\sum_{n\geq 1}\int_{u\in[(n-1)A,nA)}\widehat{f}(u,k)\mathrm{d}u,$
@@ 2.2 The Matrix exponent and Perron-Frobenius eigenvalue
$M^{(\theta)}(t)$ $=D^{(\theta)}(t)+\int_{0}^{t}D^{(\theta)}(s)CM^{(\theta)}(t-s)\mathrm{d}s+\int_{0}^{t}D^{(\theta)}(s)(Q\circ G(\theta)+E)M^{(\theta)}(t-s)\mathrm{d}s$ (2.8)
@@
$\int_{0}^{t}B^{(\theta)}M^{(\theta)}(s)\mathrm{d}s$ $=\int_{0}^{t}B^{(\theta)}D^{(\theta)}(s)\mathrm{d}s+\int_{0}^{t}B^{(\theta)}\int_{0}^{s}D^{(\theta)}(s-r)(C+Q\circ G(\theta)+E)M^{(\theta)}(r)\mathrm{d}r\mathrm{d}s$ (2.10)
@@
As mentioned in Section 1, we introduce some preliminary results on the Perron-Frobenius (PF) theory in the following Lemmas; see for example [seneta06, Theorem 1.1] or [Horn_Johnson, Theorem 8.3.4] for proofs.
@@
With the same method as in the proof of Proposition 2.13 of [Iva2011], we have the following lemma on differentiability of $\lambda(\theta)$ and $V(\theta)$.

### v2

@@ 1.4 Related works
Branching Brownian motion (BBM) and branching random walks (BRW) are canonical probabilistic models, with significant applications in statistical physics and population biology. These processes provide insights for understanding phenomena ranging from the extremes of log-correlated random fields and the structure of mean-field spin glasses to the dynamics described by diffusion-reaction equations. The classical case of a single particle type has been extensively studied. For a comprehensive treatment, we refer the reader to [54] and [11].
For branching Lévy process, there have been a recent increasing interests; see [48, 27, 29, 49], to list just a few. An extended model with infinite branching rate has been introduced by [9] and the martingale convergence has been studied in [8, 44].
The theory of branching processes has been generalized to the multitype case. Kesten and Stigum [32] established a key limit theorem for the discrete-time multitype branching processes, using the famous $L\log L$ moment condition for the reproduction law. Athreya [5] proved this result for continuous-time settings. In the context of branching Brownian motion, Ren and Yang [51] studied irreducible multitype cases, offering a probabilistic proof for the existence, uniqueness and asymptotic behaviours of the corresponding travelling wave solutions. More recently, Hou et al. [28] proved that the extremal process of an irreducible multitype branching Brownian motion converges weakly to a cluster point process. For reducible cases, two-type branching Brownian motions have been investigated in [7] and [42]. Note that, a common assumption in these works on multitype BBM has been that all particle types share the same underlying motion process. This is distinct from our setting, where the underlying processes differ among types.
A further generalization considers types in a general measurable space, which introduces significant new challenges, even without the presence of the spatial motion. There are significant new challenges. For the analysis of survival properties, we refer to [3] for an $L\log L$ condition in the model with countably many types, and to [43] for a comprehensive treatment of the case with uncountably many types.
@@ 1.5 Examples: two-type BBMs
To make connections with results in [10, 13], let us exam in detail a specific case of our model: a two-type branching Brownian motion. Particles have two types $\{1,2\}$. For $i=1,2$, a type $i$ particle undergoes a Brownian motion with drift $\mathtt{a}_{i}\in\mathbb{R}$ and variance $\sigma^{2}_{i}>0$; it branches at rate $\beta_{i}\geq 0$ into offspring of the same type at its current location, with offspring numbers distributed according to $\mu_{i}$. Furthermore, at rate $q_{1}>0$ (for type $1$) or $q_{2}>0$ (for type $2$), the particle switches to the other type without displacement. All particles evolve independently. Let $m_{i}=\sum_{k\geq 1}\mu_{i}(k)k$, $i=1,2$. Assume $\beta_{1}(m_{1}-1)+\beta_{2}(m_{2}-1)>0$ and $\sigma_{1}^{2}+\sigma_{2}^{2}>0$ to avoid degenerate cases.
@@
A connection with in [10]
This formulation of two-type branching Brownian motions encompasses the on-off branching Brownian motion models proposed by [10], where the two types are referred to as “active” and “dormant”. When $\beta_{2}=\sigma_{2}=0$, we obtain the Variant I model defined by [10, (1.16)]; when $\beta_{2}=\sigma_{1}=0$, we obtain the variant II model in [10, (1.19)].
In [10], although not explicitly stated, the last formula on page 10 has assumed $\sum k^{2}\mu_{1}(k)<\infty$. This already implies $\sum_{k}(k\log^{2}k)\mu_{j}(k)<\infty$, for $j\in\{1,2\}$. Compared to [10], our work provides several extensions:
  • For the linear speed of the leftmost particle, the result matches [10, Theorem 1.9].
  • When $\theta>\theta^{*}$, we deduce the same results on the additive martingale [10, Proposition 2.7] and travelling waves [10, Theorem 1.13].
  • When $0<\theta<\theta^{*}$, we have the same results on the additive martingale [10, Theorem 2.3 and Proposition 2.4] and travelling waves [10, Theorem 1.10].
  • When $\theta=\theta^{*}$, we include new results on martingale convergence, and also the existence and uniqueness of the travelling wave solutions, answering the open questions in [10].
A connection with [13]
The model in [13] is recovered by setting the drifts to zero ($a_{i}=0$) and requiring strictly positive branching rates ($\beta_{i}>0$ for $i=1,2$). In [13], Since the production law is given by $\mu_{1}(2)=\mu_{2}(2)=1$, it also satisfies $\mathbbm{1}_{\left\{{\beta}_{i}m_{i}>0\right\}}\sum_{k}(k\log^{2}k)\mu_{i}(k)<\infty$ for $i\in\{1,2\}$. Our results are consistent with [13]:
  • For the linear speed of the leftmost particle, the result matches the speed in [13, Theorem 1.41].
@@
  • When $0<\theta<\theta^{*}$, the $\mathcal{L}^{1}$-convergence is consistent with [13, Theorem 1.39]. We also deduce the same existence and uniqueness results in [13, Theorem 1.41].
@@ 1. Generalization to non-local branching with infinite branching rate
It is straightforward to extend our models to include non-local branching governed by a point process. We can further accommodate infinite branching rates as in [9], using similar approximation methods given there. This generalization would establish a connection with multitype growth-fragmentations [18, 17].
@@ 2. Necessary conditions for the non-triviality of the derivative martingale limit
For the single-type BBM case, (1.14) has been proven to be necessary and sufficient for the critical derivative martingale to converge to a non-trivial limit in [56]; similar results are also known as the Aïdékon–Chen condition for single-type BRWs [1, 16], and for branching Lévy processes with infinite branching rates [44]. We conjecture that an analogous statement still holds for our multitype model, namely the condition (1.14) is both necessary and sufficient. Indeed, we believe that the methods from [44] should be applicable to our model; however, a full proof would require a further study of the perpetual integral of a conditioned MAP.
@@ 3. Uniqueness of travelling waves
In Theorem 1.8, the assumption of spectrally negative jumps is used to establish uniqueness. We believe this assumption is redundant. Extending the result to processes with two-sided jumps requires a more refined analysis. We expect that the techniques from [2], which first establishes asymptotics of the travelling waves and then use these to prove uniqueness, are applicable.
@@ 4. Finer study of the leftmost position and extremal process
It would be interesting to develop a refined analysis of the leftmost position, e.g. by studying the convergence of $\min_{u\in\mathcal{N}_{t}}X_{u}(t)+\frac{\lambda(\theta^{*})}{\theta^{*}}t$ as well as the extremal process. This is closely related to the convergence of FKPP solutions to travelling waves and the precise information on the front propagation. The spine decomposition in the current work allows the methods in [54] to be adapted to the multitype setting.
@@ 5. Infinitely many types
A natural generalization is to consider more general type spaces, with countable or uncountable infinitely many types. Comparing with finitely many types, this can lead to significantly different behaviours; for example, local extinction of each type would no longer be equivalent to global extinction of the entire population [3]. This framework naturally connects to heterogeneous models of spatially dependent branching and movement. Developing such an extension would require studies on general Markov additive processes, where the underlying modulating processes are general Markov processes, as opposed to Markov chains on a finite space; see [43] for recent development.
@@ 2.1 Preliminaries on MAPs
This section provides the necessary preliminaries on Markov additive processes (MAPs) defined as in Definition 1.1. We refer to [30, Chapter 2] and [19, Appendix] for detailed discussions on this topic.
Recall from Proposition 1.2 that a MAP $(\chi,\Theta)$ is associated with a family of Lévy processes $(\chi_{i})_{i\in\mathcal{I}}$ with respective Laplace exponents $(\phi_{i})_{i\in\mathcal{I}}$, a Markov chain with intensity matrix $\Theta$ and a family of random variables $(U_{ij})_{i,j\in\mathcal{I}}$ with Laplace transform $G=(G_{ij})_{i,j\in\mathcal{I}}$. The matrix exponent of the MAP $F(\alpha)$ is given by (1.7). Since $\Theta$ is irreducible, the entries of the matrix $e^{F(\alpha)t}$ are all strictly positive. From the Perron-Frobenius (PF) theory, $F(\alpha)$ admits a PF eigenvalue which is real and larger than the real part of any other eigenvalues; see for example [53, Theorem 1.1]. Based on the PF eigenvalue, we have the following law of large numbers.
Proposition 2.1 ( [30, Propositions 2.13, 2.15 and Lemma 2.14]).
@@
The following lemma for the perpetual integral of a MAP will be used to prove Theorem 3.2. For more studies on the perpetual integrals of Lévy processes, see [6, 34].
@@
In this proof we denote by $C_{n}>0$ suitable constants for $n\geq 1$. We first treat the case when $\chi$ is non-lattice. Using [19, Theorem 27], we have
@@
Here $(R_{i,j}^{+})_{i,j\in\mathcal{I}}$ are the potential measures for the ascending ladder height process associated with $(\chi,\Theta)$, and $(R_{i,j}^{-})_{i,j\in\mathcal{I}}$ are those for the descending ladder height process; see [19, Equation (27)] for more details. When $\chi$ is lattice, then each Lévy process $\chi_{i}$ is a compound Poisson process on the same lattice $\{r\mathbb{Z}\}$ for some $r>0$, and the analysis is similar to discrete random walks. In this case, the renewal measures can be expressed via the (strong or weak) ascending ladders, analogous to the renewal measure for discrete random walks. Let $(R_{i,j}^{>})_{i,j\in\mathcal{I}}$ denote the renewal measures for the strongly ascending ladder height process of $\chi$, and let $({R}^{\leq}_{i,j})_{i,j\in\mathcal{I}}$ denote the renewal measures for the weakly descending ladder height process. Specifically, we define the renewal measures as follows. Let $T_{0}^{>}:=0$ (resp. $T_{0}^{\leq}:=0$) and define recursively for $n\geq 0$ that $T_{n+1}^{>}:=\inf\{t>T_{n}^{>}:\chi(t)>\chi(T_{n}^{>})\}$ (resp. $T_{n+1}^{\leq}:=\inf\{t>T_{n}^{\leq}:\chi(t)\leq\chi(T_{n}^{>})\}$). Let $H_{n}^{>}:=\chi(T_{n}^{>})$ (resp. $H_{n}^{\leq}:=\chi(T_{n}^{\leq})$) be the strongly ascending (resp. weakly descending) ladder heights. Then we define, for integer $m\geq 0$ and $k\leq 0$,
@@
Then by similar arguments as in the proof of [19, Theorem 27], we deduce a lattice analogue (c.f. [55, Page 209 P3] for the random walk case) of (2.2):
$\hskip 15.00002pt\mathbb{E}_{x,i}\bigg[\int_{\mathbb{R}_{+}}f(\chi(s),\Theta(s))\mathbbm{1}_{\left\{\tau_{0}>s\right\}}\mathrm{d}s\bigg]$ (2.4)
$\leq C_{1}\sum_{k\in\mathcal{I}}\sum_{j\in\mathcal{I}}\sum_{n=0}^{\infty}\sum_{m=0}^{\lfloor x/r\rfloor}R_{i,j}^{>}(\{nr\}){R}_{k,j}^{\leq}(\{mr\})f(x+(n-m)r,k)$ (2.5)
@@
It follows from the Markov renewal theorem (see e.g. [39] or [4, Section VII.4]) that they are non-negative and there exists $A>0$  [^5] Note that it holds in fact for any $A>0$ for the non-lattice case; for the lattice case, take $A$ larger than the span. such that
$R_{i,j}^{+}([nA,(n+1)A)),R_{i,j}^{-}([nA,(n+1)A)),$ (2.7)
$R_{i,j}^{>}([nA,(n+1)A)),{R}_{i,j}^{\leq}([nA,(n+1)A))\leq C_{2}A,\qquad\forall n\geq 0.$ (2.8)
@@
$\leq\sum_{n\geq 0}\widehat{f}(nA,k)C_{2}A$
$\leq C_{4}+C_{2}\sum_{n\geq 1}\int_{u\in[(n-1)A,nA)}\widehat{f}(u,k)\mathrm{d}u,$
@@ 2.2 The Matrix exponent and Perron-Frobenius eigenvalue
$\hskip 15.00002ptM^{(\theta)}(t)$ (2.11)
$=D^{(\theta)}(t)+\int_{0}^{t}D^{(\theta)}(s)CM^{(\theta)}(t-s)\mathrm{d}s+\int_{0}^{t}D^{(\theta)}(s)(Q\circ G(\theta)+E)M^{(\theta)}(t-s)\mathrm{d}s$ (2.12)
@@
$\hskip 15.00002pt\int_{0}^{t}B^{(\theta)}M^{(\theta)}(s)\mathrm{d}s$ (2.14)
$=\int_{0}^{t}B^{(\theta)}D^{(\theta)}(s)\mathrm{d}s+\int_{0}^{t}B^{(\theta)}\int_{0}^{s}D^{(\theta)}(s-r)(C+Q\circ G(\theta)+E)M^{(\theta)}(r)\mathrm{d}r\mathrm{d}s$ (2.15)
@@
As mentioned in Section 1, we introduce some preliminary results on the Perron-Frobenius (PF) theory in the following Lemmas; see for example [53, Theorem 1.1] or [26, Theorem 8.3.4] for proofs.
@@
With the same method as in the proof of Proposition 2.13 of [30], we have the following lemma on differentiability of $\lambda(\theta)$ and $V(\theta)$.